AUTOIONIZING STATES OF CALCIUM MONOFLUORIDE

\underline{XING JIANG}, JASON O. CLEVENGER, SERGEY I. PANOV AND ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

We will report experimental results from the pulsed optical-optical double resonance study of Calcium Monofluoride(CaF). These spectra are recorded via D ^{2}Σ^{+} state of CaF in the autoionization region above v^{+}=0 ionization potential with n*=15 ∼ 45. Compared to the OODR spectra of CaCl obtained previously in the same region, more ^{2}Σ^{+} and ^{2}Π states are assigned for CaF. This can be explained by the predissociation of CaCl. In this autoionization region the disappearance or shift of some members of the predicted CaF core-penetrating series reveals the existence of perturbation between certain states, which is not the case for CaCl. The OODR spectra of CaF in the high-n* region indicate that the vibrationally autoionizing states of CaF converge to v^{+} > 0 vibrational levels of CaF^{+} ion. This work completes previous studies of CaF by exploiting different intermediate states, e.g. the reverse-polarized C ^{2}Π state, to investigate the vibrational constant ω_{e} for each Rydberg series and the corresponding quantum defect derivative with respect to the internuclear distance R^{+}, which controls the autoionization process.